# Evaluating the Instruction-following Abilities of Language Models using Knowledge Tasks

### v2

@@ Abstract
LLM evaluation benchmarks have traditionally separated the testing of knowledge/reasoning capabilities from instruction following. In this work, we study the interaction between knowledge and instruction following, and observe that LLMs struggle to follow simple answer modifying instructions, and are also distracted by instructions that should have no bearing on the original knowledge task answer. We leverage existing multiple-choice answer based knowledge benchmarks and apply a set of simple instructions which include manipulating text (eg.: change case), numeric quantities (eg.: increase value, change formatting), operate on lists (eg.: sort answer candidates) and distractor instructions (eg.: change case of numeric answers).
We evaluate models at varying parameter sizes (1B-405B) from different model families and find that, surprisingly, all models report a significant drop in performance on such simple task compositions. While large-sized and frontier models report performance drops of 40-50%, in small and medium sized models the drop is severe (sometimes exceeding 80%). Our results highlight a weakness in the traditional separation of knowledge/reasoning and instruction following, and suggest that joint-study of these capabilities are important. We release our benchmark dataset, evaluation framework code, and results for future work.
Evaluating the Instruction-following Abilities of Language Models using Knowledge Tasks
Rudra Murthy IBM Research AI rmurthyv@in.ibm.com                        Praveen Venkateswaran IBM Research AI praveen.venkateswaran@ibm.com
Prince Kumar IBM Research AI prince.kumar12@ibm.com                        Danish Contractor IBM Research AI danish.contractor@ibm.com
@@ 1 Introduction
The growth of increasingly powerful large language models (LLMs) has resulted in the development of end-user applications including assistants for coding and software engineering (Ozkaya, 2023; Zhang et al., 2023; Ross et al., 2023), workflow and business automations (Grohs et al., 2023; Wornow et al., 2024), self-help assistants (Zhou et al., 2020; Shuster et al., 2022) and more. The need for highly accurate and controllable systems that follow precise instructions have led to the development of datasets and methods to improve reliability and consistency in the output for LLMs. Such methods include few-shot prompting (Gao et al., 2020; Kojima et al., 2022a), reasoning with explanations (Wei et al., 2022; Huang and Chang, 2022), checking for consistency/self-consistency (Wang et al., 2022), use of intermediate evaluators or LLMs operating as judges (Zheng et al., 2023), etc.
@@
However, despite such rapid progress and benchmark saturation we find that LLMs struggle on very simple tasks. For instance, as shown in Figure 1, given a multiple-choice question with option labels and their text, if we instruct the models to print the text associated with the correct answer instead of the answer label, we observe a significant drop ($\sim 20\%$ on average) in knowledge-task performance.
@@
From a knowledge perspective, this task is no harder than selecting (generating) the answer label. Yet, this pattern is consistent even for frontier models like GPT-4o (Appendix Figure 5). We hypothesize that this drop in performance could, in part, be due to the training process resulting in models being over-fit to certain input/output task formats, resulting in worse instruction following for other formats. However this cannot explain other observations – for instance, including additional distractor instructions that have no effect on the answering task, results in a further 5-30% drop (Section 4.2.3).
In this paper, we study this interaction of knowledge and instruction following; real-world use of LLMs typically involves both aspects, and evaluation benchmarks which specifically focus on one aspect or the other may not reflect true model performance. To keep evaluation easy and verifiable, our tasks are based on multiple-choice answers instead of long-form text generation tasks, which are harder to evaluate reliably for joint knowledge and instruction-following capabilities.
@@
We adapt commonly used knowledge and reasoning benchmarks including MMLUPro (Wang et al., 2024), MathQA Amini et al. (2019), Winogrande (Sakaguchi et al., 2021), BoolQ (Clark et al., 2019), PIQA Bisk et al. (2020) and augment them with two broad classes of instructions: (i) Instructions that are conditional on the answer to the question (ii) Instructions that are applied uniformly regardless of the answer or task. Our instructions are simple and include those that manipulate the text (eg.: change case), numeric quantities (eg: increase value, change formatting), operate on lists (eg: sort answer candidates). Our choice of using multiple-choice answers-based tasks with simple answer-modifying instructions, also allows us to automate error analysis for commonly occurring error-types and classify errors based on their origin - knowledge/reasoning errors vs instruction-following errors. Our results show that we can automatically classify approximately $80\%$ of errors.
@@
Contributions: In summary, we make the following contributions: (i) We present an evaluation framework that supports LLM-free scoring to study knowledge-conditioned instruction-following, (ii) Our experimental design enables automated error analysis (to the extent that LLMs dont have catastrophic generation failures), (iii) We create an evaluation benchmark using $5$ popular datasets, consisting of $13$ instruction types across $5$ different categories. It can be easily extended to additional tasks and instructions, (iv) We use this benchmark to evaluate $19$ models at varying parameter sizes (1B-405B) from different model families and find that, surprisingly, all models studied report a significant drop in performance on simple single-instruction task composition. While large-sized and frontier models report performance drops of 40-50%, in small and medium sized models the drop is severe (sometimes exceeding 80%). We also observe that, even within a model family, a larger parameter model may perform comparably to its smaller sibling on standard knowledge/reasoning tasks, but smaller ones drop more significantly on in our evaluation tasks. We release all code and data to enable future work.
@@ 2 Related Work
Evaluating the capabilities of large language models (LLMs) has been a significant area of research, with studies focusing on various aspects of LLM performance. Researchers have developed multiple benchmarks to assess factual knowledge Petroni et al. (2019); Roberts et al. (2020); Lin et al. (2022), logical reasoning abilities Wei et al. (2022); Zhou et al. (2023a); Saparov et al. (2023), general problem-solving capabilities Kojima et al. (2022b) and more.
Recently there have also been studies on instruction-following - for instance, FoFo Xia et al. (2024) evaluates models on format-following tasks and studies the ability of LLMs to generate outputs in existing real-world formats. In a similar vein, IFEval Zhou et al. (2023b) assesses LLMs’ ability to follow arbitrary task-specific instructions (e.g.) based on response length, casing, etc, focusing primarily on whether the instructions are followed rather than the correctness of the output for the task. InFoBench Qin et al. (2024) advances this research by introducing a metric known as the ‘Decomposed Requirements Following Ratio’ (DRFR) which is based on each aspect of an instruction that needs to be met. Along with $500$ diverse instructions and $2,250$ decomposed questions, InFoBench offers performance evaluation using OpenAI’s GPT4, across multiple constraint categories and highlights key areas where advanced LLMs can improve in complex instruction-following tasks. LLMBar (Zeng et al., 2024) is another contribution to this area, as it provides a meta-evaluation benchmark specifically designed to test an LLM evaluator’s ability to discern instruction-following outputs. The benchmark consists of $419$ manually curated pairs of outputs, where one output adheres to instructions and the other, while potentially more engaging or deceptive, does not. ComplexBench (Wen et al., 2024) assesses the ability of LLMs to follow arbitrary task specific instructions, while Meowpass (He et al., 2024) presents a benchmark to evaluate complex instruction following by dynamically composing multiple IFEval style instructions, and consists of $1500$ diverse prompt and instruction pairs. Another study introduces ManyIFEval (Harada et al., 2025), a benchmark assessing LLMs’ ability to follow multiple instructions, revealing a “curse of instructions” where accuracy declines as instruction count increases. Li et al. (2024) propose a method to evaluate instruction following ability via verbalizer manipulation. Specifically, they modify the classification task labels with different verbalizers which may or may not be semantically relevant to the task. They observe that all models fail to follow instructions when they instruct the model to flip the labels (unnatural setting).
@@ 3.2 Instruction Categories
3.3 Metrics and Output Post-processing
[…]
3.4 Automated classification of errors
@@ 3.5.1 Benchmark ranking
Instruction Category Score (IC Score): We compute the micro-average exact-match scores for every instance per instruction category and then compute the arithmetic mean.
Knowledge Task Subset Score: (KTS Score): We compute the micro-average exact-match scores for every instance per knowledge-task, and then compute the arithmetic mean.
Exact Match Score on ‘Instructions with no-effect’ ($\mu_{EM}^{{}^{\prime}}$): We compute the micro-average of all instruction instances in the benchmark that have no effect on the original knowledge-task answers (i.e.) ‘distractors’.
@@ 4.2.3 Effect of distractor instructions
In the appendix Figures 6, 7, 8, we report details of how different model families (Llama, Qwen, Phi) are affected by distractors, at different scales. In summary, we find that all models are distracted by instructions that require reversing and casing text,555partially applicable (can only reverse non-single digit numeric answers) or completely inapplicable reporting a drop of 55-75%. Even frontier models (Appendix Figure 9) show a drop of 35%. However, for the other distractor instructions, we find that Llama models are relatively robust, showing a drop of $5-10\%$, while the Phi and Qwen family of models are more susceptible, where we observe drops of 20-30%. Frontier models surprisingly also show a drop of 10-15%. While model failures in the presence of distractors have been studied before (Shi et al., 2023; Feng et al., 2024), to the best of our knowledge this is the first work to study them in an instruction-following setting.
@@ 4.2.4 Automated Error Classification
Figure 4 shows the error analysis for two different model families - Llama and Qwen2.5.666We include error analysis plots for other model families in the appendix. We report the average exact match score as well as the errors.
@@
In addition, we observe that the Qwen2.5 family of models in almost always better than the Llama family of models when it comes to numeric manipulation instructions while the Llama family of models has a slightly higher performance on string manipulation tasks (Appendix Figures 40 and 40). A similar trend is observed on the unclassified set where Llama’s instruction-following failures appear to be more severe. The Phi family of models (Appendix Figure 34) appears to perform worse than both the Qwen and Llama family of models. Frontier models (Appendix Figure 38) demonstrate lower knowledge errors as expected, but still make many instruction following errors.
@@ 5 Discussion & Conclusion
In this work, we took a novel approach to studying instruction-following by grounding instructions on existing knowledge tasks. Our approach has the advantage of being easily extendable for new instruction types and domains, while also enabling LLM-free evaluations with some degree of automated error analysis. We demonstrated that not only do models fail to follow simple instructions (e.g.) printing the answer text instead of the label, but their performance drops further when compound but simple, instructions are included. Even when instructions that should have no effect on the knowledge-tasks are used, models at all scales report a drop in performance, though the extent of deterioration varies. As models are increasingly being viewed as agents and assistants, it is crucial that models have better guarantees of following user instructions. As our work demonstrates, there is a lot of scope for improvement and we hope the community finds it helpful in improving the current state-of-the-art.
6 Limitations
@@ Appendix A Appendix
We begin by including details of the knowledge tasks used. We then describe our data creation process and how we automatically classify errors in section 3.4. We list all instructions with example input, ground truth, and expected instruction output in B.4. We report results on the Full Benchmark in Table 5. The detailed statistics of Full and Lite Benchmark are presented in B.5. Section C.1 presents the comparison between model’s performance on print correct answer and print correct answer labels tasks on the Lite Benchmark. Section C.3 presents performance of different models for each instruction category in comparison with its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Table 5: Performance of the Medium, Large and Frontier Models on our Full Benchmark - models ranked in order of performance using the average score (higher is better).
@@ B.3 Knowledge and Reasoning Tasks
(i) MMLUPro (Wang et al., 2024): MMLUPro extends the MMLU dataset to make it more challenging by a) increasing the number of options from four to ten and b) increasing problem difficulty by focusing on more reasoning oriented problems. We consider all $14$ subjects in the MMLUPro benchmark. We cap the maximum number of samples for each subject to be $150$ samples.
(ii) MathQA (Amini et al., 2019): MathQA dataset consists of math word problems presented as Multiple-Choice-Questions (MCQs). Given a math question and four options, the model has to select the correct answer.
@@
(iii) BoolQ (Clark et al., 2019): BoolQ is a boolean question-answering dataset. Given a passage and a boolean question around the passage, the model has to select either True or False.
(iv) PIQA (Bisk et al., 2020): Physical Interaction: Question Answering (PIQA) involves answering questions that involve commonsense reasoning around physical objects. Given a question and two options, the model has to select the most plausible option.
(v) Winogrande (Sakaguchi et al., 2021): Winogrande involves a fill-in-the-blank task with binary options, the model has to select the correct option for a given sentence. The task involves reasoning for pronoun resolution.
@@ B.4 Prompts
B.5 Additional Benchmark Statistics

### v3

# KCIF: Knowledge-Conditioned Instruction Following

@@ Abstract
LLM evaluation benchmarks have traditionally separated the testing of knowledge/reasoning capabilities from instruction following. In this work, we study the interaction between knowledge and instruction following, and observe that LLMs struggle to follow simple answer modifying instructions, and are also distracted by instructions that should have no bearing on the original knowledge task answer. We leverage existing multiple-choice answer based knowledge benchmarks and apply a set of simple instructions which include manipulating text (eg.: change case), numeric quantities (eg.: increase value, change formatting), operate on lists (eg.: sort answer candidates) and distractor instructions (eg.: change case of numeric answers). We evaluate models at varying parameter sizes (1B-405B) from different model families and find that, surprisingly, all models report a significant drop in performance on such simple task compositions. While large-sized and frontier models report performance drops of 40-50%, in small and medium sized models the drop is severe (sometimes exceeding 80%). Our results highlight a limitation in the traditional separation of knowledge/reasoning and instruction following, and suggest that joint-study of these capabilities are important. We release our benchmark dataset, evaluation framework code, and results for future work.
@@ 1 Introduction
The need for highly accurate and controllable LLM-powered systems that follow precise instructions have led to the development of datasets and methods to improve reliability and consistency in the output for LLMs. Such methods include few-shot prompting (Gao et al., 2020; Kojima et al., 2022b), reasoning with explanations (Wei et al., 2022; Huang and Chang, 2022), checking for consistency/self-consistency (Wang et al., 2022), use of intermediate evaluators or LLMs operating as judges (Zheng et al., 2023), and more.
@@
However, despite such rapid progress and ‘benchmark saturation’ LLMs can often struggle on very simple tasks. For instance, as shown in Figure 1, given a multiple-choice question with option labels and their text, if we instruct the models to print the text associated with the correct answer instead of the answer label, we observe a significant drop ($\sim 20\%$ on average) in knowledge-task performance.
@@
From a user’s perspective, this task is no harder than selecting (generating) the answer label. Yet, this pattern is consistent even for frontier models like GPT-4o (Appendix Figure 5) and such patterns of interaction where keys or values are referred to LLM instructions are fairly common in real-world LLM usage – for instance, in manipulating lists (eg: “replace the third bulleted item with ‘x”’ ), operating on semi-structured objects (eg: extracting the value of a JSON field and use in a subsequent operation), etc. While this example of a drop in performance when printing the answer-text instead of the answer-label could, be due to the training process resulting in models being over-fit to certain input/output task formats, or due to issues associated with label binding (Xue et al., 2024), there are other patterns of failure as well, – for instance, including additional distractor instructions that have no effect on the answering task, results in a further 5-30% drop (Section 4.2.3).
@@
In this paper, we study this interaction of knowledge and instruction following; real-world use of LLMs typically involves both aspects, and ‘nested’ instructions where there is conditional branching of instructions based on intermediate steps is a common usage pattern of LLMs. However, evaluating these reliably can be hard and can require the use of complex LLM-judges (He et al., 2024). Thus, to keep evaluation easy and verifiable, our tasks are based on multiple-choice answers instead of long-form text generation task, We adapt commonly used knowledge and reasoning benchmarks including MMLUPro (Wang et al., 2024), MathQA Amini et al. (2019), Winogrande (Sakaguchi et al., 2021), BoolQ (Clark et al., 2019), PIQA Bisk et al. (2020) and augment them with two broad classes of instructions: (i) Instructions that are conditional on the answer to the question (ii) Instructions that are applied uniformly regardless of the answer or task. Our instructions are simple and include those that manipulate the text (eg.: change case), numeric quantities (eg: increase value, change formatting), operate on lists (eg: sort answer candidates). Our choice of using multiple-choice answers-based tasks with simple answer-modifying instructions, also allows us to automate error analysis for commonly occurring error-types and classify errors based on their origin - knowledge/reasoning errors vs instruction-following errors. Our results show that we can automatically classify approximately $80\%$ of errors.
@@
Contributions: In summary, we make the following contributions: (i) We present an evaluation framework that supports LLM-free scoring to study knowledge-conditioned instruction-following, (ii) Our novel experimental design enables automated error analysis (to the extent that LLMs dont have catastrophic generation failures), (iii) We create an evaluation benchmark using $5$ popular datasets, consisting of $13$ instruction types across $5$ different categories. It can be easily extended to additional tasks and instructions, (iv) We use this benchmark to evaluate $19$ models at varying parameter sizes (1B-405B) from different model families and find that, all models studied report a significant drop in performance on simple single-instruction task composition. While large-sized and frontier models report performance drops of 40-50%, in small and medium sized models the drop is severe (sometimes exceeding 80%). We also observe that, even within a model family, a larger parameter model may perform comparably to its smaller sibling on standard knowledge/reasoning tasks, but smaller ones drop more significantly on in our evaluation tasks. We release code [^1] https://github.com/IBM/KCIF and data [^2] https://ibm.biz/KCIFData to support evaluation, enable task extensions and future work.
@@ 2 Related Work
Evaluating the capabilities of large language models (LLMs) has been a significant area of research, with studies focusing on various aspects of LLM performance. Researchers have developed multiple benchmarks to assess factual knowledge Petroni et al. (2019); Roberts et al. (2020); Lin et al. (2022), logical reasoning abilities Wei et al. (2022); Zhou et al. (2023a); Saparov et al. (2023), general problem-solving capabilities Kojima et al. (2022a) and more.
Recently there have also been studies on instruction-following - for instance, FoFo Xia et al. (2024) evaluates models on format-following tasks and studies the ability of LLMs to generate outputs in existing real-world formats. In a similar vein, IFEval Zhou et al. (2023b) assesses LLMs’ ability to follow arbitrary task-specific instructions (e.g.) based on response length, casing, etc, focusing primarily on whether the instructions are followed rather than the correctness of the output for the task. InFoBench Qin et al. (2024) advances this research by introducing a metric known as the ‘Decomposed Requirements Following Ratio’ (DRFR) which is based on each aspect of an instruction that needs to be met. Along with $500$ diverse instructions and $2,250$ decomposed questions, InFoBench offers performance evaluation using OpenAI’s GPT4, across multiple constraint categories and highlights key areas where advanced LLMs can improve in complex instruction-following tasks. LLMBar (Zeng et al., 2024) is another contribution to this area, as it provides a meta-evaluation benchmark specifically designed to test an LLM evaluator’s ability to discern instruction-following outputs. The benchmark consists of $419$ manually curated pairs of outputs, where one output adheres to instructions and the other, while potentially more engaging or deceptive, does not. ComplexBench (Wen et al., 2024) and FollowBench (Jiang et al., 2024) aim to assess the ability of LLMs to follow arbitrary task specific instructions, while Meowpass (He et al., 2024) presents a benchmark to evaluate complex instruction following by dynamically composing multiple IFEval style instructions, and consists of $1500$ diverse prompt and instruction pairs. Another study introduces ManyIFEval (Harada et al., 2025), a benchmark assessing LLMs’ ability to follow multiple instructions, revealing a “curse of instructions” where accuracy declines as instruction count increases. Li et al. (2024) propose a method to evaluate instruction following ability via verbalizer manipulation. Specifically, they modify the classification task labels with different verbalizers which may or may not be semantically relevant to the task. They observe that all models fail to follow instructions when they instruct the model to flip the labels (unnatural setting).
@@ 3.2 Instruction Categories
3.3 Metrics and Automated Error Classification
We report the model performance as a form of exact match (EM) where we perform basic string parsing (removing beginning and ending whitespaces, quotations, etc.) and compare the model prediction to the expected output for the applied instruction. See Appendix Section B.5 for more details.
@@ 3.4.1 Benchmark ranking
Instruction Category Score (IC Score): To ensure models perform well across instruction categories, we also compute the micro-average exact-match scores for every instance per instruction category and then compute the arithmetic mean.
Knowledge Task Subset Score: (KTS Score): To ensure model capabilities generalize across different knowledge tasks, we compute the micro-average exact-match scores for every instance per knowledge-task, and then compute the arithmetic mean.
Exact Match Score on ‘Instructions with no-effect’ ($\mu_{EM}^{{}^{\prime}}$): Lastly, we compute the micro-average of all instruction instances in the benchmark that have no effect on the original knowledge-task answers (i.e.) ‘distractors’.
@@ 4.2.3 Effect of distractor instructions
In the appendix Figures 6, 7, 8, we report details of how different model families (Llama, Qwen, Phi) are affected by distractors, at different scales. In summary, we find that all models are distracted by instructions that require reversing and casing text,666partially applicable (can only reverse non-single digit numeric answers) or completely inapplicable reporting a drop of 55-75%. Even frontier models (Appendix Figure 9) show a drop of 35%. However, for the other distractor instructions, we find that Llama models are relatively robust, showing a drop of $5-10\%$, while the Phi and Qwen family of models are more susceptible, where we observe drops of 20-30%. Frontier models surprisingly also show a drop of 10-15%. While model failures in the presence of distractors have been studied before (Shi et al., 2023; Feng et al., 2024), to the best of our knowledge this is the first work to study them in an instruction-following setting.
@@ 4.2.4 Automated Error Classification
Figure 4 shows the error analysis for two different model families - Llama and Qwen2.5.777We include error analysis plots for other model families in the appendix. We report the average exact match score as well as the errors.
@@
In addition, we observe that the Qwen2.5 family of models in almost always better than the Llama family of models when it comes to numeric manipulation instructions while the Llama family of models has a slightly higher performance on string manipulation tasks (Appendix Figures 40 and 40). A similar trend is observed on the unclassified set where Llama’s instruction-following failures appear to be more severe. The Phi family of models (Appendix Figure 34) appears to perform worse than both the Qwen and Llama family of models. Frontier models (Appendix Figure 38) demonstrate lower knowledge errors as expected, but still make many instruction following errors. We include a qualitative example of each error category in Appendix Section B.6.
@@ 5 Discussion & Conclusion
In this work, we took a novel approach to studying instruction-following by grounding instructions on existing knowledge tasks. Our approach has the advantage of being easily extendable for new instruction types and domains, while also enabling LLM-free evaluations with some degree of automated error analysis. We demonstrated that not only do models fail to follow simple instructions (e.g.) printing the answer text instead of the label, but their performance drops further when compound but simple, instructions are included. Even when instructions that should have no effect on the knowledge-tasks are used, models at all scales report a drop in performance, though the extent of deterioration varies. As models are increasingly being viewed as agents and assistants, it is crucial that models have better guarantees of following user instructions. As our work demonstrates, there is a lot of scope for improvement and we hope the community finds our framework helpful in improving the current state-of-the-art.
@@ Appendix A Appendix
We begin by including details of the knowledge tasks used. We then describe our data creation process and how we automatically classify errors in section 3.3. We list all instructions with example input, ground truth, and expected instruction output in B.4. We report results on the Full Benchmark in Table 5. The detailed statistics of Full and Lite Benchmark are presented in B.7. Section D.1 presents the comparison between model’s performance on print correct answer and print correct answer labels tasks on the Lite Benchmark. Section D.3 presents performance of different models for each instruction category in comparison with its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Table 5: Performance of the Small, Medium, and Large Models on our Full Benchmark - models ranked in order of performance using the average score (higher is better).
@@ B.3 Knowledge and Reasoning Tasks
(i) MMLUPro [Wang et al., 2024]: MMLUPro extends the MMLU dataset to make it more challenging by a) increasing the number of options from four to ten and b) increasing problem difficulty by focusing on more reasoning oriented problems. We consider all $14$ subjects in the MMLUPro benchmark. We cap the maximum number of samples for each subject to be $150$ samples.
(ii) MathQA [Amini et al., 2019]: MathQA dataset consists of math word problems presented as Multiple-Choice-Questions (MCQs). Given a math question and four options, the model has to select the correct answer.
@@
(iii) BoolQ [Clark et al., 2019]: BoolQ is a boolean question-answering dataset. Given a passage and a boolean question around the passage, the model has to select either True or False.
(iv) PIQA [Bisk et al., 2020]: Physical Interaction: Question Answering (PIQA) involves answering questions that involve commonsense reasoning around physical objects. Given a question and two options, the model has to select the most plausible option.
(v) Winogrande [Sakaguchi et al., 2021]: Winogrande involves a fill-in-the-blank task with binary options, the model has to select the correct option for a given sentence. The task involves reasoning for pronoun resolution.
@@ B.4 Prompts
B.5 Metrics and Output Post-processing
[…]
B.6 Qualitative examples of error types
B.6.1 Type 1: Instruction following error
Model: Llama-3.1-8B-Instruct
Input Task (Append correct answer with a string):
Given a sentence with a blank and 2 options namely ’A’ and ’B’, print the text associated with the option label that answers the question correctly. Do not print the option label. Without including any additional characters, also append the following text string at the end of the output:’Date’. Do not add any special characters including quotations, spaces, new lines etc in the response. Think step by step and in the end, finish your response with ’Response:\$RESPONSE’ where \$RESPONSE (without quotes) is the final output expected.
Sentence: Samantha strictly followed their diet while Sarah ate what they felt like, and _ ate chicken and broccoli all week.
Options: A. Samantha B. Sarah
Ground Truth:
SamanthaDate
Output:
Date
Response:Samantha
B.6.2 Type 2: Knowledge Error
Model: Qwen 2.5-72B-instruct
Input Task (Print the correct answer text:
Given a question about biology and 10 options: A, B, C, D, E, F, G, H, I, J as candidate answers, print the text associated with the option label that answers the question correctly. Do not print the option label. Think step by step and in the end, finish your response with ’Response:\$RESPONSE’ where \$RESPONSE (without quotes) is the final output expected.
Question: Nymphaea odorata, a water lily, is an aquatic plant whose leaves and flowers float on the surface of the water. Which of the following characteristics would be expected for leaves of this plant?
A. Stomata on the bottom surface with no water-conducting (xylem) cells
B. stomata on the bottom surface and relatively few water-conducting (xylem) cells
C. Stomata on the top surface with no water-conducting (xylem) cells
D. Stomata evenly distributed on both surfaces and relatively many water-conducting (xylem) cells
E. stomata on the top surface and relatively few water-conducting (xylem) cells
F. No stomata and relatively many water-conducting (xylem) cells
G. stomata on the top surface and relatively many water-conducting (xylem) cells
H. No stomata and relatively few water-conducting (xylem) cells
I. Stomata evenly distributed on both surfaces and relatively few water-conducting (xylem) cells
J. stomata on the bottom surface and relatively many water-conducting (xylem) cells
Ground Truth:
stomata on the top surface and relatively few water-conducting (xylem) cells
Output:
To determine the correct answer, let’s analyze the characteristics of the leaves of Nymphaea odorata, a water lily.
1. **Stomata Location**:
- Stomata are essential for gas exchange, which is crucial for photosynthesis and transpiration. In aquatic plants, especially those with floating leaves, stomata are typically located on the upper surface of the leaves. This is because the upper surface is exposed to the air, allowing for efficient gas exchange, while the lower surface is in contact with water, which would impede gas exchange.
2. **Water-Conducting Cells (Xylem)**:
- Xylem is responsible for transporting water and minerals from the roots to the leaves. In aquatic plants, the availability of water is not a limiting factor, so the need for extensive xylem tissue is reduced. However, some xylem is still necessary to transport minerals and support the plant.
Given these points, we can eliminate options that do not align with these characteristics:
- Options A, B, and J, which suggest stomata on the bottom surface, are incorrect because the bottom surface is in contact with water, which would not allow for efficient gas exchange.
- Options C, E, and G, which suggest no stomata or stomata on the top surface, are more plausible. However, we need to consider the xylem content.
- Options F and H, which suggest no stomata, are incorrect because stomata are necessary for gas exchange.
- Options D and I, which suggest stomata evenly distributed on both surfaces, are less likely because the bottom surface is in contact with water, making stomata on the bottom less effective.
The most plausible option is:
- **G. stomata on the top surface and relatively many water-conducting (xylem) cells**
This option aligns with the need for stomata on the top surface for gas exchange and the presence of some xylem for mineral transport and support.
Response: stomata on the top surface and relatively many water-conducting (xylem) cells’
B.6.3 Type 3: Both Knowledge and Instruction following errors
Model: Gemma-2-9b-it
Input Task - Capitalize correct answer:
Given a sentence with a blank and 2 options namely ’A’ and ’B’, capitalize the text associated with the optional label that answers the question correctly. Do not print the option label. Think step by step and in the end, finish your response with ’Response:\$RESPONSE’ where \$RESPONSE (without quotes) is the final output expected.
Sentence: John could not hear his alarm clock when he was sleeping with a headphone on his head because the \_ is closer. Options: A. headphone B. clock
Ground truth:
HEADPHONE
Model Output:
Let’s analyze the sentence step by step:
- The sentence is about John not being able to hear his alarm clock.
- The blank is where we need to insert either headphone or clock’.
- The sentence states that John is wearing headphones on his head.
- Therefore, the alarm clock would be closer to his head than the headphones.
- This means the alarm clock would be the closer object.
- The correct option to fill in the blank is clock’.
Now, let’s capitalize the text associated with the correct option.
Response:clock
B.6.4 Type 4: Unclassified Errors
Model: Qwen 2.5-3B-instruct
Input Task (Print the correct answer text:
Given a mathematical question and 5 options namely ’a’, ’b’, ’c’, ’d’, and, ’e’, as candidate answers, print the text associated with the option label that answers the question correctly. Do not print the option label. Without including any additional characters, also append the following text string at the end of the output:’xQbkTIPMZ7wlBJCx’. Do not add any special characters including quotations, spaces, new lines etc in the response. Think step by step and in the end, finish your response with ’Response:\$RESPONSE’ where \$RESPONSE (without quotes) is the final output expected.
Question: two friends c and d leave point c and point d simultaneously and travel towards point d and point c on the same route at their respective constant speeds . they meet along the route and immediately proceed to their respective destinations in 32 minutes and 50 minutes respectively . how long will d take to cover the entire journey between point d and point c ?
Options:
a. 90
b. 80
c. 75
d. 60
e. 65
Output:
To solve this problem, let’s denote the distance between point C and point D as \(D\), the speed of friend C as \(v_C\), and the speed of friend D as \(v_D\). Let’s also denote the time they meet as \(t\) minutes.
When they meet, the sum of the distances they have traveled is equal to the total distance \(D\). So, we have:
\[ v_C t + v_D t = D \]
\[ t (v_C + v_D) = D \]
[…]
After meeting, friend C takes 32 minutes to reach point D, and friend D takes 50 minutes to reach point C. This means:
[…]
Simplifying these equations:
[…]
….. continued incoherent reasoning and no final response provided
B.7 Additional Benchmark Statistics